OpenStudy (anonymous):

how do you factorise 3(x^2-20)-31x? pls help!!!!

OpenStudy (callisto):

3(x^2-20)-31x = 3x^2 -31x - 60 = (x-12) (3x+5)

OpenStudy (anonymous):

thanks!!

OpenStudy (callisto):

welcome